two cars leave the same place at the same time, one going east and the other west. Both are going at an average speed of 22mph. In how many hours will they be 220 miles apart?
Answer by lwsshak3(11628)

two cars leave the same place at the same time, one going east and the other west. Both are going at an average speed of 22mph. In how many hours will they be 220 miles apart?
***
cars are moving away from each other at a speed of 22+22=44 mph
travel time=distance/speed=220/44=5
In how many hours will they be 220 miles apart? 5